A particle is constrained to move along the cardioid r = 2 ( 1 + cose ˈθ ) where r is in meters. Its position vector has constant angular speed of 2 rad/s counterclockwise. At what rate is the particle changing its speed when ˈθ=60 ?
1
Expert's answer
2018-08-15T09:01:08-0400
r=2·(1+cos⁡〖60 ̊)=〗 2·(1+1/2)=3 metres
v=r·ω=3·2=6 m/s